OpenStudy (anonymous):

Find cot θ if csc θ = square root5/2and tan θ > 0

OpenStudy (ybarrap):

What do you know about \(\csc \theta\) and \(\cot\theta\) with respect to a right triangle?

OpenStudy (anonymous):

Could you use the pythagorean identities?

OpenStudy (ybarrap):

yes

OpenStudy (ybarrap):

Can you draw a picture that shows csc θ = square root5/2? Let's start with that.

OpenStudy (anonymous):

 Sorry Idk if I'm doing this right.

OpenStudy (ybarrap):

Let's start with that.  \(\Large\csc\theta={1\over \sin\theta}\). So \(\Large \sin\theta={1\over \csc\theta}\) . How do we find \(\Large \cos\theta\)?